# Merging Facts, Crafting Fallacies: Evaluating the Contradictory Nature of Aggregated Factual Claims in Long-Form Generations

### v2

@@ Abstract
Long-form generations from large language models (LLMs) contain a mix of factual and non-factual claims, making evaluating factuality difficult. To evaluate factual precision of long-form generations in a more fine-grained way, prior works propose to decompose long-form generations into multiple verifiable facts and verify those facts independently. The factuality of the generation is the proportion of verifiable facts among all the facts. Such methods assume that combining factual claims forms a factual paragraph. This paper shows that the assumption can be violated. We show that LLMs can generate paragraphs that contain verifiable facts, but the facts are combined to form a non-factual paragraph due to entity ambiguity. We further reveal that existing factual precision metrics, including FActScore and citation recall, cannot properly evaluate the factuality of these non-factual paragraphs. To address this, we introduce an enhanced metric, D-FActScore, specifically designed for content with ambiguous entities. We evaluate the D-FActScores of people biographies generated by retrieval-augmented LLMs. We show that D-FActScore can better assess the factuality of paragraphs with entity ambiguity than FActScore. We also find that four widely used open-source LLMs tend to mix information of distinct entities to form non-factual paragraphs.
@@ 1 Introduction
Figure 1: Output of an LLM when prompted to generate a biography for Dick Hanley. While the paragraph is misleading and non-factual, all the facts in the paragraph can be supported by the Wikipedia of Dick Hanley (Swimmer)  or Dick Hanley (American FootBall) , yielding 100% FActScore. D-FActScore groups atomic facts that appear to refer to the same individual based on the paragraph (Figure 1(c)), finds an entity that best matches that individual from the knowledge source (Figure 1(d)), and only uses the information of that entity to verify the facts in that group .
The correctness of the factuality metrics mentioned above relies on the assumption that "as long as each claim is factual, the combination of those claims is factual". This paper shows that this assumption can be violated due to entity ambiguity in the generation. Figure 1 shows such a case, where a paragraph mixes the information of two entities with the same name, but readers without prior or external knowledge will think all the information refers to the same individual. Since the paragraph is not written in a way that allows readers to understand the content factually, the paragraph is non-factual. However, since each claim in the generation is supported by Wikipedia, the FActScore of the above misleading paragraph is 100% by definition.
We emphasize the significance of this evaluation problem by showing that LLMs can easily generate this kind of non-factual paragraph. We prompt LLMs to generate a biography for a target name using retrieval-augmented generation (Guu et al., 2020; Lewis et al., 2020), where the target name is shared by multiple entities, and the retrieved documents may also include the Wikipedia pages of multiple entities. We find that Llama-13b/70b-chat (Touvron et al., 2023), Tulu-13b-dpo (Ivison et al., 2023), and Vicuna-7b (Chiang et al., 2023) tend to mix the information of multiple entities in the same biography, and a reader without prior knowledge cannot tell that the information corresponds to different entities. Non-factual as these paragraphs are, existing factuality metrics cannot correctly assess the factuality in this case.
@@
We recruit humans to annotate the D-FActScores of people biographies generated with Llama-13b-chat, Tulu-v2-13b-dpo, and ChatGPT (OpenAI, 2022). We show that Llama and Tulu mix the information of multiple entities in a non-factual way, and D-FActScore can identify those non-factual generations. We also propose a pipeline to estimate D-FActScore automatically and show that our automatic evaluation pipeline estimates D-FActScore to a reasonable degree.
@@ 2.1 Collecting Names with Ambiguity
To generate biographies with mixed information of multiple entities with the same name, we collect names with ambiguity from Wikipedia disambiguation pages using the Wikipedia API. A disambiguation page lists all the Wikipedia pages of the entities with the same name. We randomly select 500 ambiguous names and call this collection AmbigBio in our paper. We use target name to refer to a name in AmbigBio. A target name corresponds to multiple entities in Wikipedia.
@@ 2.2 Prompting LLMs to Generate Biographies
We use 2-shot demonstrations to prompt the LLM. The demonstration is similar to the prompt in Table 1, and the <answer> is replaced with a paragraph written by the authors. We categorize the demonstrations into two types:
(1) With name ambiguity: The names in the demonstrations are ambiguous names in Wikipedia; the retrieved passages include Wikipedia pages of different entities with the same name. The <answer> only contains the information of one of the entities with that name.
(2) Without name ambiguity: The name in the demonstration is unambiguous (there is no disambiguation page for that name in Wikipedia). The retrieved results contain the passages from that entity’s Wikipedia page and possibly some unrelated passages. The <answer> is the bio of that entity.
@@ 2.4 Categorizing LLM-Generated Paragraphs
We categorize the generated paragraphs based on the number of distinct entities and the number of disambiguable biographies, defined as follows. The below definitions and categorization are better understood with the examples in Table 2.
@@
Type LLM-generated paragraph
Metrics
One bio one entity [Bio1] [Ent1] John Stewart, a Northern Irish politician born in 1983, has served as a Member of the Northern Ireland Assembly (MLA) for East Antrim since 2017. Stewart is a Unionist and previously worked in sales for Robinson’s Shoemakers, a family business. Before joining politics, he attended Carrickfergus Grammar School and earned a Bachelor of Arts (BA) degree from Cardiff University.
FS: 92% D-FS: 92% Num bio: 1 Num ent: 1
One bio $N$ entities [Bio1] [Ent1] John Hemming (born January 5, 1935) is a British historian and explorer who has made significant contributions to the study and understanding of the Incas and indigenous peoples of the Amazon Basin. Born in Vancouver to a British journalist mother, Alice Hemming OBE, and a father who had served in the First World War and wanted to ensure his son’s birth in North America, John Hemming was raised in England from the age of two months. [Ent2] Hemming went on to study at Clapham Grammar School before [Ent3] earning a degree in Arithmetic and Theoretical, Atomic and Nuclear Physics at Magdalen College, Oxford.
FS: 100% D-FS: 71% Num bio: 1 Num ent: 3
$N$ bios $N$ entities [Bio1] [Ent1] Joseph F. Smith, the sixth President of The Church of Jesus Christ of Latter-day Saints, was born in 1838, and was the nephew of Joseph Smith, the founder of the Latter Day Saint movement. Additionally, Smith, who was the last president to have personally known the church’s founder, led the LDS Church. [Bio2] [Ent2] Joseph F. Smith, an American politician from Pennsylvania, was also born in 1920 and served in the United States House of Representatives. After a decorated military career, Smith was elected to represent Pennsylvania during the Ninety-seventh United States Congress.
FS: 87% D-FS: 87% Num bio: 2 Num ent: 2
@@
(2) One biography and $N$ entities: The generated output mixes the information of different entities with the target name but does not provide sufficient disambiguation information. A typical reader without prior knowledge will consider the whole paragraph a single bio of one individual.
(3) $N$ biographies and $N$ entities: The paragraph contains information about multiple entities and provides enough context for readers without prior knowledge to disambiguate different entities in the paragraph. Each biography describes one of the entities with the target name.
@@ 3 Existing Factual Precision Metrics
Min et al. (2023a) show that FActScore can be obtained using an automatic pipeline: they instruct GPT3.5 (text-davinci-003) to decompose a long-form generation into atomic facts and use another $\text{LM}_{\text{EVAL}}$ to verify each fact. They propose three different types of $\text{LM}_{\text{EVAL}}$ to verify an atomic fact: (1) No-context LM: prompt an $\text{LM}_{\text{EVAL}}$ using ’<atomic fact> True or False?’ (2) Retrieve $\to$ LM: retrieve $k$ passages from Wikipedia, construct the prompt by concatenating the retrieved passages, the atomic fact, and ’True or False?’, and prompt $\text{LM}_{\text{EVAL}}$ to answer. (3) Nonparametric Probability (NP): this is calculated by the average token likelihood of the atomic fact using a nonparametric masked LM (Min et al., 2023b). The atomic fact is considered True if the average token likelihood exceeds a pre-defined threshold. Min et al. (2023a) show that ensembling NP and Retrieve $\to$ LM can approximate the FActScore obtained by human annotation.
@@ 4.1 Definition
For each group of atomic fact $\mathcal{A}_{y,i}$, we use entity linking to find an entity $e_{i}^{*}$ in the knowledge source $\mathcal{C}$ that best matches the facts in $\mathcal{A}_{y,i}$. Let the subset of the knowledge source related to $e_{i}^{*}$ denoted by $C_{i}^{*}$. For all atomic facts $a\in\mathcal{A}_{y,i}$, they will be verified using $C_{i}^{*}$ instead of using the whole $C$. The D-FActScore of $y$ is defined as follows:
@@ 4.3 Human Evaluation Results
Table 3: Human and automatic evaluations are conducted on the same set of paragraphs. FS: FActScore, D-FS: D-FActScore, # bio: average number of separable biographies in one paragraph, # ent.: average number of distinct entities in one paragraph, chat-13b: Llama-13b-chat.
@@ 5.3 Different LLMs and Demonstrations
After showing that D-FActScore can be estimated using automatic evaluation, we use automatic evaluation to evaluate the D-FActScores of paragraphs generated using all 500 names in AmbigBio by five LLMs and two types of demonstrations. We have the following findings.
Demonstrations with name ambiguity do not make the outputs more factual.   In Table 4, we compare the results obtained with two types of demonstrations: names with and without ambiguity. Recall that in the demonstration with name ambiguity, the target name is a name with ambiguity, and the top-5 retrieved documents contain passages from different entities with the same name. Since the demonstration <answer> only contains a biography that includes the information of one entity, we hope the LLM can know how to handle target names with ambiguity better. However, we do not see a higher D-FActScore when using demonstrations with name ambiguity.
@@ 6 Related Work
We discuss the related works in the appendix due to the limited space. If the paper is accepted, we plan to move this section to the main content using the additional page in the camera-ready version.
@@
Evaluating the factuality of texts is an active subfield in NLP. Some prior works formulate factuality evaluation as an uncertainty estimation problem, and they consider generation that models are less confident to be less factual and more likely to be hallucination (Liu et al., 2022; Zhang et al., 2023b; Manakul et al., 2023). Many recent works focus on evaluating the precision of the citation when generating texts with citations and use QA or NLI models to determine if the statements in the generation can be supported by the cited documents (Rashkin et al., 2023; Liu et al., 2023; Gao et al., 2023; Yue et al., 2023). Our work is largely based on FActScore (Min et al., 2023a), and we improve FActScore by resolving its inability to handle entity ambiguity.
@@
Unfaithfulness due to incorrect or incomplete coreference is a well-known problem in summarization (Pagnoni et al., 2021; Zhang et al., 2023a). The problem of combining facts into non-factual paragraphs identified in our paper is related to incorrect coreference but is different since incorrect coreference stems from ambiguous anaphors instead of ambiguous entities. The unfaithfulness due to coreference ambiguity is very hard to evaluate using automatic evaluation metrics, as automatic evaluation metrics have low correlations with human evaluation results (Pagnoni et al., 2021). D-FActScore, which can be estimated by automatic evaluation, can properly evaluate the non-factual contents stemming from entity ambiguity or incorrect coreference, which is one of our contributions compared to prior works.
@@
Our work uses an ambiguous name to prompt the LLM to generate a biography. This is highly related to ambiguous question answering (Min et al., 2020; Stelmakh et al., 2022), where a question can have multiple answers based on how the question is interpreted. Min et al. (2020) estimates that about 23% of the questions in AmbigNQ dataset are due to entity ambiguity, which is the focus of our work. Most prior works answer an open-domain ambiguous question by first generating disambiguated questions and generating an answer for each disambiguated questions (Gao et al., 2021; Min et al., 2021; Shao and Huang, 2022; Kim et al., 2023). Some works do not disambiguate the ambiguous question and generate the answer directly (Stelmakh et al., 2022; Gao et al., 2023). We do not perform the disambiguation step in our work because we want to know if the LLMs can perform well when we do not explicitly disambiguate the target entity.
@@ 7 Conclusion and Discussion
Our findings underscore LLMs’ difficulties in generating accurate content when retrieving from Wikipedia with entity ambiguity. The scenario in our paper is simplified, given that Wikipedia is not the sole information source about entities. Even using high-quality retrieved content from Wikipedia, open-source models struggle to differentiate between entities accurately. In more realistic situations where LLMs draw from a broader and more nuanced content pool on the Web, and the named entities used to query LLMs may not appear in the training data, distinguishing between different entities becomes even more challenging. Overcoming this issue is vital for the reliability of retrieval-augmented LLMs.
@@
Despite the limitations listed above, we believe that our paper significantly contributes to identifying a realistic evaluation problem that has never been studied and proposing a simple solution.
We do not see specific risks or harm of our paper.

### v4

@@ Abstract
Long-form generations from large language models (LLMs) can contain a mix of factual and non-factual claims, making evaluating factuality difficult. Prior works evaluate the factuality of a long paragraph by decomposing it into multiple facts, verifying those facts independently, and aggregating the results. Such methods assume that combining factual claims forms a factual paragraph. The above assumption can be violated: we show that strong open-source models like Llama-chat can generate paragraphs that contain verifiable facts, but the facts are combined into a non-factual paragraph due to entity ambiguity. We further reveal that existing factuality metrics, including FActScore and citation recall, cannot properly evaluate these non-factual paragraphs and overestimate their factuality. To address this, we introduce an enhanced metric, D-FActScore, specifically designed for content with ambiguous entities. We evaluate the D-FActScores of people biographies generated by retrieval-augmented LLMs. We show that D-FActScore can better assess the factuality of paragraphs with entity ambiguity than FActScore. We also find that four widely used open-source LLMs tend to mix information of distinct entities to form non-factual paragraphs, making their D-FActScore much lower than FActScore by over 10%.
@@ 1 Introduction
Figure 1: Output of Llama-13b-chat when prompted to generate a biography for Dick Hanley. While the paragraph is misleading and non-factual, all the facts in the paragraph can be supported by the Wikipedia of Dick Hanley (Swimmer)  or Dick Hanley (American FootBall) , yielding 100% FActScore. D-FActScore groups atomic facts that appear to refer to the same individual based on the paragraph (Figure 1(c)), finds an entity that best matches that individual from the knowledge source (Figure 1(d)), and only uses the information of that entity to verify the facts in that group .
The correctness of the previous factuality metrics relies on the assumption that "as long as each claim is factual, the combination of those claims is factual." This paper shows that this assumption can be violated due to entity ambiguity in the generation. Figure 1 shows such a case, where Llama-2-13b-chat says that Dick Hanley was born in 1936 and died in 1970, where the birth date is true for Dick Hanley (Swimmer) and the time of death is true for Dick Hanley (American Football), but readers will think all the information refers to the same individual. Since the paragraph is not written in a way that allows readers to understand the content factually, the paragraph is non-factual. However, since each claim in the generation is supported by Wikipedia, the FActScore of the above misleading paragraph is 100% by definition.
We emphasize the significance of this evaluation problem by showing that LLMs can easily generate this kind of non-factual paragraph. We collect 500 ambiguous names from Wikipedia to form a dataset, AmbigBio, where each name corresponds to multiple entities in Wikipedia. We prompt LLMs to generate a biography for a name in AmbigBio using retrieval-augmented generation (Guu et al., 2020; Lewis et al., 2020), where the target name is shared by multiple entities, and the retrieved documents may also include the Wikipedia pages of multiple entities. We find that Llama-13b/70b-chat (Touvron et al., 2023), Tulu-13b-dpo (Ivison et al., 2023), and Vicuna-7b (Chiang et al., 2023) tend to mix the information of multiple entities in the same biography, and a reader without prior knowledge cannot tell that the information corresponds to different entities. Non-factual as these paragraphs are, existing factuality metrics cannot correctly assess the factuality in this case.
@@
We recruit humans to annotate the D-FActScores of people biographies generated with Llama-13b-chat, Tulu-v2-13b-dpo, and ChatGPT (OpenAI, 2022). We show that Llama and Tulu mix the information of multiple entities in a non-factual way. Compared to FActScore, D-FActScore correctly identifies those non-factual generations and gives a D-FActScore 8% to 16% lower than FActScore. We also propose a pipeline to estimate D-FActScore automatically and show that our automatic pipeline estimates D-FActScore to a reasonable degree.
@@
    We collect 500 ambiguous names into a new dataset, AmbigBio, and each name corresponds to multiple entities in Wikipedia.
  3. 3.
@@ 2.1 Collecting Names with Ambiguity
To generate biographies with mixed information of multiple entities with the same name, we collect names with ambiguity from Wikipedia disambiguation pages using the Wikipedia API. A disambiguation page lists all the Wikipedia pages of the entities with the same name. We randomly select 500 ambiguous names and call this collection AmbigBio in our paper. We use target name to refer to a name in AmbigBio. On average, a target name corresponds to 4.6 entities in Wikipedia.
@@ 2.2 Prompting LLMs to Generate Biographies
We use 2-shot demonstrations to prompt the LLM. The demonstration is similar to the prompt in Table 1, and the <answer> is replaced with a paragraph written by the authors.
@@ 2.4 Categorizing LLM-Generated Paragraphs
We categorize the generated paragraphs based on the number of distinct entities and the number of disambiguable biographies, defined as follows. The below definitions and categorization are better understood with the examples in Table 2.4.
@@
{tblr}
colspec = X[m, 6em] X[m, 35em] X[m, 5em], row1 = font= Type LLM-generated paragraph Metrics
One bio one entity [Bio1] [Ent1] John Stewart, a Northern Irish politician born in 1983, has served as a Member of the Northern Ireland Assembly (MLA) for East Antrim since 2017. Stewart is a Unionist and previously worked in sales for Robinson’s Shoemakers, a family business. Before joining politics, he attended Carrickfergus Grammar School and earned a Bachelor of Arts (BA) degree from Cardiff University. FS: 92% D-FS: 92% Num bio: 1 Num ent: 1
One bio $N>1$ entities [Bio1] [Ent1] John Hemming (born January 5, 1935) is a British historian and explorer who has made significant contributions to the study and understanding of the Incas and indigenous peoples of the Amazon Basin. Born in Vancouver to a British journalist mother, Alice Hemming OBE, and a father who had served in the First World War and wanted to ensure his son’s birth in North America, John Hemming was raised in England from the age of two months. [Ent2] Hemming went on to study at Clapham Grammar School before [Ent3] earning a degree in Arithmetic and Theoretical, Atomic and Nuclear Physics at Magdalen College, Oxford. FS: 100% D-FS: 71% Num bio: 1 Num ent: 3
$N>1$ bios $N>1$ entities [Bio1] [Ent1] Joseph F. Smith, the sixth President of The Church of Jesus Christ of Latter-day Saints, was born in 1838, and was the nephew of Joseph Smith, the founder of the Latter Day Saint movement. Additionally, Smith, who was the last president to have personally known the church’s founder, led the LDS Church. [Bio2] [Ent2] Joseph F. Smith, an American politician from Pennsylvania, was also born in 1920 and served in the United States House of Representatives. After a decorated military career, Smith was elected to represent Pennsylvania during the Ninety-seventh United States Congress. FS: 87% D-FS: 87% Num bio: 2 Num ent: 2
@@
(2) One biography and $N>1$ entities: The generated output mixes the information of different entities with the target name but does not provide sufficient disambiguation information. A typical reader without prior knowledge will consider the whole paragraph a single bio of one individual.
(3) $N>1$ biographies and $N>1$ entities: The paragraph contains information about multiple entities and provides enough context for readers without prior knowledge to disambiguate different entities in the paragraph. Each biography describes one of the entities with the target name.
@@ 3 Existing Factual Precision Metrics
Min et al. (2023a) show that FActScore can be obtained using an automatic pipeline: they instruct GPT3.5 (text-davinci-003) to decompose a long-form generation into atomic facts and use another $\text{LM}_{\text{EVAL}}$ to verify each fact. They propose three different types of $\text{LM}_{\text{EVAL}}$ to verify an atomic fact: (1) No-context LM: prompt an $\text{LM}_{\text{EVAL}}$ using ’<atomic fact> True or False?’ (2) Retrieve $\to$ LM: retrieve $k$ passages from Wikipedia, construct the prompt by concatenating the retrieved passages, the atomic fact, and ’True or False?’, and prompt $\text{LM}_{\text{EVAL}}$ to answer. (3) Nonparametric Probability (NP): an atomic fact is factual if the average token likelihood of the fact calculated using a nonparametric masked LM (Min et al., 2023b) exceeds a pre-defined threshold.
@@ 4.1 Definition
For each group of atomic facts $\mathcal{A}_{y,i}$, we use entity linking to find an entity $e_{i}^{*}$ in the knowledge source $\mathcal{C}$ that best matches the facts in $\mathcal{A}_{y,i}$. Let the subset of the knowledge source related to $e_{i}^{*}$ denoted by $C_{i}^{*}$. For all atomic facts $a\in\mathcal{A}_{y,i}$, they will be verified using $C_{i}^{*}$ instead of using the whole $C$. The D-FActScore of $y$ is defined as follows:
@@ 4.3 Human Evaluation Results
Table 3: Human and automatic evaluations of FActScore and D-FActScore conducted on the same set of paragraphs. FS: FActScore, D-FS: D-FActScore, # bio: average number of separable biographies in one paragraph, # ent.: average number of distinct entities in one paragraph, chat-13b: Llama-13b-chat.
@@ 5.2 Human VS Automatic Evaluation
Per-paragraph D-FActScore of human and automatic evaluation results show a high correlation.   We calculate the per-paragraph Pearson correlation coefficient $r$ between D-FActScore obtained with human evaluation and automatic evaluation, and we also calculate $r$ between the number of bios obtained by human and automatic evaluation. The results shown in Table 5 in the Appendix reveal that the Person’s $r$ of D-FActScore between human and automatic evaluation is moderate (ChatGPT) to strong (Llama and Tulu). If we consider the D-FActScore of the paragraphs generated by all three models together and calculate Pearson’s r, the correlation coefficient is 0.75, which is a strong correlation. We also find that for all three models, the number of bios obtained using human and automatic evaluation are all very strongly correlated (Pearson’s $r\geq 0.8$). The above experiment results further prove the effectiveness of the automatic evaluation pipeline.
@@ 5.3 Different LLMs and Demonstrations
After showing that D-FActScore can be estimated using automatic evaluation, we use automatic evaluation to evaluate the D-FActScores of paragraphs generated using all 500 names in AmbigBio by five LLMs and two types of demonstrations: with and without name ambiguity.
(1) With name ambiguity: The names in the demonstrations are ambiguous names in Wikipedia; the retrieved passages include Wikipedia pages of different entities with the same name. This is the default setting in the previous experiments.
(2) Without name ambiguity: The name in the demonstration is unambiguous (there is no disambiguation page for that name in Wikipedia). The retrieved results contain the passages from that entity’s Wikipedia page and possibly some unrelated passages.
The <answer> in the demonstrations with and without name ambiguity only contains the information of one entity with the target name. We have the following findings.
Demonstrations with name ambiguity do not make the outputs more factual.   In Table 4, we compare the results of demonstrations with and without ambiguity. In the with ambiguity demonstration, the demonstration <answer> only contains a biography that includes the information of one entity, and we hope the LLM can know how to handle target names with ambiguity better. However, we do not see a higher D-FActScore when using demonstrations with name ambiguity.
@@ 6 Related Work
Evaluating the factuality of texts is an active subfield in NLP. Some prior works formulate factuality evaluation as an uncertainty estimation problem and consider generation that models are less confident to be less factual (Liu et al., 2022; Zhang et al., 2023b; Manakul et al., 2023). Many recent works focus on evaluating the precision of the citation when generating texts with citations and use QA or NLI models to determine if the statements in the generation can be supported by the cited documents (Rashkin et al., 2023; Liu et al., 2023; Gao et al., 2023; Yue et al., 2023). Our work is largely based on FActScore (Min et al., 2023a), and we improve FActScore by resolving its inability to handle entity ambiguity.
@@
Unfaithfulness due to incorrect or incomplete coreference is a well-known problem in summarization (Pagnoni et al., 2021; Zhang et al., 2023a). The problem of combining facts into non-factual paragraphs identified in our paper is related to incorrect coreference but is different since incorrect coreference stems from ambiguous anaphors instead of ambiguous entities. Unfaithfulness due to coreference ambiguity is very hard to evaluate using automatic evaluation metrics, as automatic evaluation metrics have low correlations with human evaluation results Pagnoni et al. (2021). D-FActScore, which can be estimated by automatic evaluation, can properly evaluate the non-factual contents stemming from entity ambiguity.
@@
Our work uses an ambiguous name to prompt the LLM to generate a biography. This is highly related to ambiguous question answering (Min et al., 2020; Stelmakh et al., 2022), where a question can have multiple answers based on how the question is interpreted. Min et al. (2020) estimates that about 23% of the questions in AmbigNQ dataset are due to entity ambiguity, which is the focus of our work. Most prior works answer an open-domain ambiguous question by first generating disambiguated questions and generating an answer for each disambiguated questions (Gao et al., 2021; Min et al., 2021; Shao and Huang, 2022; Kim et al., 2023). Some works do not disambiguate the ambiguous question and generate the answer directly (Stelmakh et al., 2022; Gao et al., 2023). We do not perform the disambiguation step because we want to know if the LLMs can perform well when we do not explicitly disambiguate the target entity.
@@ 7 Conclusion and Discussion
Our findings underscore LLMs’ difficulties in generating accurate content when retrieving from Wikipedia with entity ambiguity. The scenario in our paper is simplified, given that Wikipedia is not the sole information source about entities. Even using high-quality contents from Wikipedia, open-source models struggle to differentiate between entities accurately. In more realistic situations where LLMs draw from a broader and more nuanced content pool on the Web, and the named entities used to query LLMs may not appear in the training data, distinguishing between different entities becomes even more challenging. Overcoming this issue is vital for the reliability of LLMs with RAG.
@@
Using Wikipedia for RAG and fact-checking
We use Wikipedia for RAG and fact-checking. This is a very common protocol in the literature of RAG research since Wikipedia is the most commonly used knowledge source. For example, Self-RAG (Asai et al., 2023) also retrieves documents from Wikipedia to generate biographies and uses Wikipedia to evaluate the factuality of the generated biographies. It is important to explore the case when the RAG does not use Wikipedia for generation, and we will leave extending the data to other sources as future works.
Fact-checking beyond Wikipedia
In our paper, we evaluate D-FActScore by relying on Wikipedia of the source for fact-checking. This is a common protocol in fact-checking, which uses a structured knowledge source like Wikipedia or other databases to verify the facts. In the case of verifying facts using unstructured knowledge sources, more effort may be required to disambiguate entities in the knowledge source. To properly calculate D-FActScore, we need to be able to split the information in the knowledge source into the information of distinct entities. That is, we need to disambiguate the articles in the knowledge source, for example, verifying the biography of a CS PhD student who is not in Wikipedia and using the contents on the Web for verification. Let’s say we find a LinkedIn page and some semantic scholar pages with the same name as our target entity. To disambiguate the entities, we need to match semantic scholar pages with LinkedIn profiles. This can be done by various methods, including but not limited to (1) checking if the LinkedIn profile already is linked to a semantic scholar page, or (2) checking if the LinkedIn profile is linked to the person’s homepage and checking if one can find a semantic scholar link on that page. This can all be done automatically with tool-augmented and retrieval-augmented LLMs. However, the above process may be challenging when the knowledge source is more complicated and provides too little information for disambiguation. We leave verifying facts using the contents on the Web in future work.
@@
We do not see specific risks or harm in our paper.
Acknowledgements
We thank the reviewers for providing detailed feedback and actionable suggestions, which helped us strengthen our paper. We also thank the senior committee members for monitoring the reviewing process. We thank Sewon Min for providing valuable, actionable, concrete feedback and insights on the paper. Cheng-Han Chiang is supported by a Google PhD Fellowship and a Ph.D. scholarship program by Delta Electronics. We thank the National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.